Trace the conicoid represented by x^2+2z^2°y. Also describe its section by planes x=c, for all c belongs to R.
1
Expert's answer
2020-03-16T13:24:42-0400

Trace the conicoid represented by "x^2+2z^2=y" . Also describe its sections by

the planes "x=c,\forall c\in \R."

It follows from the conicoid equation "y=x^2+2z^2" that solution exists only for "y\gt 0" .

For "y\lt 0," there is no solution with "x,z\in \R." The form of conicoid is displayed on fig.1 .

For "y\gt0," conicoid sections with planes "y=c\in\R\gt0" have form of ellipses with the formul

"a\frac{x^2}{c}+\frac{z^2}{c/2}=1" (fig.2). The center of the ellipse is located on the axis "Y". The small axis is

directed along "Z" and has a length of "b=\sqrt{c/2}." The large axis of the ellipses is direced

along the "X" axis an has a length of "a=\sqrt{c}." The formula for this section in the new notations

take on a canonical form of ellipses "\frac{x^2}{a^2}+\frac{z^2}{b^2}=1". The conicoid body takes on value within

"-a\lt x\lt a;-b\lt z\lt b."

The section by "x=c\in \R" has the equation "y=c^2+2z^2" which is a parabola with an axis

os symmetry parallel ith axis "Y." The canonical form of the parabola is

"(y-c^2)=\frac{z^2}{2\cdot p}" , where "p=1\/4," and "y=c^2" is its vertex. The vertex moves out of plane

"XZ(y=0)" as c increases. All the section are similar to each other with focus distance of

parabola "F=\frac{p}{2}=1/8." Due to symmetry about "YZ" plane for "c\lt0" we get exactly

the same cross sections.

fig.1


fig.2